OpenStudy (rootbeer003):

Which equation would best help solve the following problem: Brett kicks a field goal with an initial vertical velocity of 42 meters per second. How long will it take the football to hit the ground? –4.9t^2 – 42t = 0 –16t^2+ 42t = 0 –4.9t^2 + 42t = 0 16t^2 + 42t = 0

OpenStudy (anonymous):

C

OpenStudy (anonymous):

No it's b

OpenStudy (rootbeer003):

guys

OpenStudy (rootbeer003):

i need truth

OpenStudy (anonymous):

It's definetely b

OpenStudy (anonymous):

Daniel 56k answered 2 of my questions earlier and they were wrong don't listen to him

OpenStudy (rootbeer003):

how do u know its b

OpenStudy (anonymous):

i am sure it was option 'c' is right.

OpenStudy (anonymous):

Because I looked it up and did research

OpenStudy (ffasinger):

yeah i think it is c too

OpenStudy (anonymous):

Lies

OpenStudy (anonymous):

You'll thank me later rootbeer

OpenStudy (ffasinger):

dude @Edude23 chill

OpenStudy (baconboy73):

Think about it. How fast is gravity when related to time? What about when it is related to displacement? Also think about the football's initial velocity. Is it in the same direction as gravity, or opposite?

OpenStudy (rootbeer003):

need the right answer

OpenStudy (anonymous):

i can show you, it is simple 2d-motion question, use equation\[s=ut+(1/2)at ^{2}\]there distance covered 's' is 0, u is initial velocity, a is gravitational accelration, (take negative, i.e. -g), and you will got your answer.

OpenStudy (rootbeer003):

Which equation would best help solve the following problem: DeSean throws a ball up with an initial vertical velocity of 30 feet per second from a platform that is at ground level. How long will the ball be in the air? 16t^2 + 30t = 0 –16t^2+ 30t = 0 –16t^2 – 30t = 0 –4.9t^2 – 30t = 0

OpenStudy (rootbeer003):

what about this?

OpenStudy (anonymous):

It's either a b c or d or nah

OpenStudy (rootbeer003):

edude get off my questions

OpenStudy (ffasinger):

that was uncalled for

OpenStudy (ffasinger):

i believe it is B

OpenStudy (anonymous):

Yes sir

OpenStudy (rootbeer003):

same

OpenStudy (anonymous):

I'm sorry I bothered you

OpenStudy (ffasinger):

you were just being rude chill bro

OpenStudy (anonymous):

Alright

OpenStudy (rootbeer003):

Martina hits a volleyball ball upward with an initial velocity of 13 meters/sec. When she hits the ball, it is one meter above the ground. How long will it take the ball to hit the ground? –4.9t^2 + 13t + 1 = 0 –4.9t^2 + 13t = 1 4.9t^2 + 13t + 1 = 0 –16t^2+ 13t + 1 = 0

OpenStudy (rootbeer003):

and this??

OpenStudy (anonymous):

I'm sorry I bothered you

OpenStudy (anonymous):

Yes sir

OpenStudy (ffasinger):

i think d

OpenStudy (ffasinger):

anything elese rootbeer

OpenStudy (rootbeer003):

ty gurl

OpenStudy (ffasinger):

welcome any time!

OpenStudy (anonymous):

i think it is...

OpenStudy (anonymous):

definitely not D!!

OpenStudy (rootbeer003):

yea that was wrong LOLOLO

OpenStudy (rootbeer003):

BUT THE REST WERE RIGHT

OpenStudy (anonymous):

D would be for the measure in feet not meters

OpenStudy (ffasinger):

sorry i was not positive i was baseing it of the first 1 lol

OpenStudy (rootbeer003):

its all good